# SAEMark: Multi-bit LLM Watermarking with Inference-Time Scaling

### v1

@@ Abstract
Watermarking LLM-generated text is critical for content attribution and misinformation prevention. However, existing methods compromise text quality, require white-box model access and logit manipulation—limitations that exclude API-based models and multilingual scenarios. We propose SAEMark, a general framework for post-hoc multi-bit watermarking that embeds personalized messages solely via inference-time, feature-based rejection sampling without altering model logits or requiring training. Our approach operates on deterministic features extracted from generated text, selecting outputs whose feature statistics align with key-derived targets. This framework naturally generalizes across languages and domains while preserving text quality through sampling LLM outputs instead of modifying. We provide theoretical guarantees relating watermark success probability and compute budget that hold for any suitable feature extractor; empirically, we demonstrate the framework’s effectiveness using Sparse Autoencoders (SAEs), achieving superior detection accuracy and text quality. Experiments across 4 datasets show SAEMark’s consistent performance, with 99.7% F1 on English and strong multi-bit detection accuracy. SAEMark establishes a new paradigm for scalable watermarking that works out-of-the-box with closed-source LLMs while enabling content attribution.  [^1] We open-source code and data at: https://zhuohaoyu.github.io/SAEMark
@@ 1 Introduction
Large language models (LLMs) have revolutionized text generation across domains, from creative writing to code synthesis (Brown et al., 2020; Guo et al., 2024). However, their ability to produce human-quality text at scale raises serious concerns about misinformation, copyright infringement, and content laundering. As these models become ubiquitous, reliably attributing AI-generated content becomes critical for accountability and trust.
Watermarking—embedding detectable signatures into generated text—offers a promising solution, but existing approaches face a fundamental tradeoff. They must preserve text quality while enabling reliable detection, operate across languages and domains, and scale to distinguish between many users or sources. Most critically, they must work with real-world deployment constraints where model providers offer only API access without exposing internal parameters.
The challenge becomes even more complex for multi-bit watermarking. Beyond simply detecting AI-generated text, the goal is to encode and recover a specific message $m\in\{0,1\}^{b}$—such as a user identifier for personalized attribution. This enables answering not just “is this AI-generated?” but “which specific user or system generated this text?” Such fine-grained attribution is essential for large-scale deployment where accountability matters.
Existing watermarking methods struggle with these requirements. Token-level approaches like KGW (Kirchenbauer et al., 2023) and EXP (Aaronson & Kirchner, 2022) require direct access to model logits, excluding API-based deployment, and can degrade text quality through probability manipulation. Syntactic methods (Hou et al., 2023) fail to generalize across languages, while specialized approaches (Lee et al., 2024) work well in narrow domains but break down when applied more broadly. Even recent black-box methods (Bahri & Wieting, 2024; Chang et al., 2024) rely on surface-level statistics or require auxiliary models, limiting their robustness and scalability.
@@
This approach works by operating on meaningful units of text—sentences for natural language, functions for code. For each unit, we extract deterministic features that capture semantic properties, compute a scalar statistic, and normalize it to behave predictably across different texts. Using the watermark key, we derive target values for each position. During generation, we sample multiple candidates from the LLM and select the one whose feature statistic is closest to the target, ensuring the final sequence encodes the desired message.
The elegance lies in what we don’t change: no model weights, no logit manipulation, no token modifications. Every selected text segment is a natural LLM output, preserving quality while enabling attribution. The approach works with any LLM through API calls, generalizes across languages and domains, and provides theoretical guarantees on watermark success that scale predictably with computational budget.
@@
Our contributions span theory and practice. We develop a general framework for watermarking through feature-guided selection that works with any feature extractor and any language model API. We provide theoretical guarantees that explain how watermark success scales with computational resources and text length, independent of the specific features used. Finally, we demonstrate a practical instantiation using Sparse Autoencoders that achieves superior detection accuracy and text quality across English, Chinese, and code, encoding more information per unit length than existing multi-bit approaches.
Figure 1: An overview of SAEMark.
2 Preliminaries
2.1 Related Work
LLM watermarking is a technique to embed special patterns into the output of LLMs, and has traditionally been used to identify LLM generated text from human-written text (Jawahar et al., 2020). Different from post-hoc detection methods (Zellers et al., 2019) that analyze statistical patterns in existing text, language model watermarking aims to embed detectable signatures during generation (Kirchenbauer et al., 2023).
@@ 2.1 Related Work
Existing approaches exhibit several limitations. Many methods compromise generation quality through direct manipulation of token probabilities (Kirchenbauer et al., 2023) or syntactic modifications (Atallah et al., 2002). The challenge of language and domain generalization remains largely unaddressed, with current techniques primarily optimized for English and struggling with multilingual content or specialized domains like code (Lee et al., 2024). Notably, PersonaMark (Zhang et al., 2024b) represents early attempts at personalized watermarking, but its reliance on English-specific syntactic patterns and closed-source implementation makes scalability and cross-lingual capability difficult to verify. Recently, more multi-bit watermarking methods have been proposed to embed multiple bits of information into generated text (Lau et al., 2024; Zhang et al., 2024a; Yoo et al., 2023a; Guan et al., 2024; Yoo et al., 2023b; Qu et al., 2024), primarily by extending single-bit watermarking that manipulates logits during generation; these methods inherit the limitations of single-bit designs.
@@
Complementary black-box watermarks avoid white-box logit access by using post-hoc selection or rewriting (Chang et al., 2024). However, they typically operate on surface statistics or introduce auxiliary model dependencies and do not directly address multi-bit message embedding at scale. Our framework differs by performing inference-time selection among naturally generated candidates using deterministic feature statistics computed over domain-appropriate units. This enables extractor-agnostic analysis and multilingual, domain-agnostic multi-bit watermarking without modifying model logits.
2.2 Sparse Autoencoders
@@
Sparse Autoencoders (SAEs) are pre-trained interpretability tools that decompose LLM activations into human-understandable features (Bricken et al., 2023). For a given base model $\mathcal{M}$ and layer $l$, an SAE processes hidden states $\mathbf{h}_{t}$ at position $t$ as:
@@
This training produces features that correspond to interpretable concepts (Bricken et al., 2023; Templeton et al., 2024). For instance, SAEs applied to Gemma 2B (Team et al., 2024)’s final transformer layer, GemmaScope, reveal features like ”Python function definitions” or ”Concept related to color blue” (Lieberum et al., 2024).
Our watermarking leverages three key properties of pre-trained SAE features. First, layer-specific patterns capture distinct behaviors from different model layers. Second, multilingual activation allows the same features to fire for equivalent concepts across languages. Third, sparsity enables efficient analysis through few active features per token. These properties support language-agnostic statistics via masked feature aggregations:
@@
where $\mathbf{m}$ filters background features that fire ubiquitously regardless of content (e.g., punctuation-associated features). The summary $\phi(\mathbf{y})$ provides a deterministic statistic used by our generation and detection procedures (Sec 3).
@@ 2.3 Task Definition
Given a base LLM $\mathcal{M}$, watermark key $k\in\mathcal{K}$, input $\mathbf{x}$, and message $m\in\{0,1\}^{b}$, the algorithm produces $\mathbf{y}$ by post-hoc selection over $\mathcal{M}$’s outputs (black-box/API-compatible; no access to logits or parameters; cf. black-box watermarking (Bahri & Wieting, 2024; Chang et al., 2024)):
@@
The scheme targets three properties: key privacy (deriving $k$ from watermarked outputs is hard), verifier-held detectability (any party holding $k$ can verify), and collusion resistance (multiple keys should not facilitate removal or forgery).
Threat model and scope.
Our focus is attribution without storing sensitive content. This work does not claim cryptographic unforgeability when keys are known; preventing adversarial forgeries under black-box constraints is an important direction for security-focused follow-ups.
3 Methodology
@@ 3.1 General Framework for Feature-based Watermarking
where $M$ denotes the number of textual units in the sequence and $O_{\min}=0.95$ ensures that at least 95% of target values fall within the observed range.
These two filters aim to eliminate spurious matches: the range similarity filter prevents matching sequences with fundamentally different statistical properties, while the overlap rate filter ensures meaningful correspondence between target and observed values. Only after passing both alignment checks do we apply Student’s t-test for statistical significance. The key with the highest significance score is returned if it passes the threshold; otherwise, we classify the text as unwatermarked.
Input: Prompt $c$, key $k$, LLM $G$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, units $M$, attempts $K$, candidates NN, alignment thresholds
Output: Watermarked text $x^{*}$
for $attempt\leftarrow 1$ to $K$ do
 for $i\leftarrow 1$ to $M$ do
    $\mathcal{X}\leftarrow\text{GenerateCandidates}(G,x^{*},N)$
   end for
 if CheckAlignment($\{\tau_{i}\},\{z_{i}\}$then
    return $x^{*}$
   end if
end for
return $x^{*}$
Algorithm 1 Watermark generation
Input: Text $x$, candidate keys $\mathcal{K}$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, alignment thresholds, significance level $\alpha$
Output: Detection result $d\in\mathcal{K}\cup\{\emptyset\}$
$\mathcal{D}\leftarrow\emptyset$
foreach $k_{i}\in\mathcal{K}$ do
 $\{\tau_{j}\}\leftarrow\text{TargetsFromKey}(k_{i},|\{z_{j}\}|)$
 if CheckAlignment($\{\tau_{j}\},\{z_{j}\}$then
    $t,p\leftarrow\text{StudentTTest}(\{z_{j}\},\{\tau_{j}\})$
    if $t>t_{\alpha/2}\land p<\alpha$ then
       $\mathcal{D}\leftarrow\mathcal{D}\cup\{(k_{i},t)\}$
      end if
   end if
end foreach
Algorithm 2 Watermark detection
Figure 2: Pseudocode for SAEMark: generation and detection.
@@ 3.2 Theoretical Analysis and Guarantees
Worst-case analysis and tight bounds.
@@
where $\Phi$ denotes the standard normal CDF. This simplifies to $p_{\min}=\Phi(2(1+k)+k\mu/\sigma)-\Phi(2(1-k)-k\mu/\sigma)$.
@@ 3.3 Sparse Autoencoder Instantiation
What concrete feature extractor should we use? We need statistics that are deterministic, semantically meaningful, and statistically regular. Sparse autoencoders—interpretability tools designed to understand language model internals—provide an ideal solution. They decompose language representations into interpretable semantic components (”technical writing,” ”narrative voice,” ”mathematical reasoning”) that exhibit distinctly different activation patterns across generations. By applying the sparse autoencoder to a separate ”anchor” model, our approach remains compatible with any target language model, including API services, while extracting the discriminative yet predictable statistics our framework requires.
@@
Two practical optimizations bridge the gap between our theoretical framework and robust empirical performance. First, the CheckAlignment algorithm’s eliminate spurious statistical matches that would otherwise compromise detection accuracy. Second, background feature masking ensures FCS calculations focus on discriminative semantic patterns rather than ubiquitous surface features. We precompute a mask excluding ”background” SAE features, like those related to punctuation or basic grammar, to focus on discriminative semantic patterns. With empirically observed parameters $\mu=0.142$, $\sigma=0.029$, and tolerance $\epsilon=0.1$, our bounds yield concrete success probabilities: $N=50$ achieves $99\%$ per-unit success, $N=20$ maintains $85.32\%$, and $N=10$ achieves $61\%$. Since each generation involves multiple units, overall success rates significantly exceed these per-unit bounds. Modern inference engines support parallel generation of $N$ candidates simultaneously, making the approach practically efficient despite these extra compute overhead. We have extensive ablations and empirical results in the following experiments.
@@ 4.1 Experimental Setup
C4 (2020) LCSTS (2015) MBPP (2021) PandaLM (2023b)
@@
Task Type
Completion
Summarization
Code Gen.
Inst. Following
@@
Setup.
We evaluate on 4 diverse datasets as shown in Table 1. Following common practice in prior work, we report Accuracy, Recall, F1 at 1% FPR. For text quality, we report win-rates of pairwise comparison on PandaLM judged by GPT-4o in our main results, and average pointwise scores on BIGGen-Bench judged by their officially released judge model as an alternative text-quality experiment. We use implementation for single-bit watermarks from MarkLLM (Pan et al., 2024) toolkit and Waterfall (Lau et al., 2024) as it’s the current best open-source training-free multi-bit watermark similar to our setting. Full details of baselines in Appendix C.
@@ 4.2 Watermarking Accuracy and Text Quality
Method C4 (English, 2020) LCSTS (Chinese, 2015) MBPP (Code, 2021) PandaLM (Instruction,  2023b)
@@
KGW (2023) 99.2 99.6 99.2 99.1 98.8 99.1 65.4 31.9 48.0 41.5 89.9 80.4 88.8
EXP (2022) 99.5 99.6 99.5 99.3 99.4 99.3 57.8 16.7 28.4 23.2 79.3 59.4 74.2
UPV (2023) 86.0 72.0 83.7 90.5 91.0 90.5 51.6 3.1 6.0 36.0 54.0 8.0 14.8
Unigram (2023) 98.8 98.6 98.8 98.2 97.0 98.2 65.4 31.9 48.0 35.3 53.3 7.2 13.4
DIP (2023) 96.0 92.6 95.9 97.7 96.2 97.7 60.7 22.6 36.5 36.5 81.5 63.8 77.5
Unbiased (2023b) 96.7 94.4 96.6 97.8 96.4 97.8 64.0 29.2 44.8 40.2 74.3 49.3 65.7
SynthID (2024) 98.2 97.2 98.2 97.6 96.2 97.6 62.5 26.1 41.0 36.0 81.2 63.0 77.0
SWEET (2024) 99.6 99.6 99.6 50.0 0.0 0.0 72.4 45.9 62.4 47.2 87.7 76.8 86.2
@@
Beyond accuracy, Table 2 shows SAEMark achieves the highest quality score (67.6%) on PandaLM as judged by GPT-4o pairwise comparisons. To study how this conclusion generalizes across different backbone models, since backbone performance is the most significant factor affecting text quality, we conduct additional evaluation on BIGGen-Bench comparing against both watermarked baselines and unwatermarked text.
Table 3: Text quality evaluation on BIGGen-Bench. Scores are on a 5-point Likert scale (higher is better), judged by the officially released BIGGen-Bench judge model (Kim et al., 2024).
@@
Table 3 confirms SAEMark consistently achieves the highest quality among watermarking methods across three backbone models. This quality advantage stems from a fundamental difference in approach: rather than manipulating logits or applying external rewriting to obtain watermarked text, we simply select among naturally generated candidates. This post-hoc selection guarantees that every watermarked segment remains an unmodified LLM generation from the backbone model itself, ensuring text quality stays bounded by the its own capabilities.

### v2

# SAEMark: Steering Personalized Multilingual LLM Watermarks with Sparse Autoencoders

@@ Abstract
Watermarking LLM-generated text is critical for content attribution and misinformation prevention, yet existing methods compromise text quality and require white-box model access with logit manipulation or training, which exclude API-based models and multilingual scenarios. We propose SAEMark, an inference-time framework for multi-bit watermarking that embeds personalized information through feature-based rejection sampling, fundamentally different from logit-based or rewriting-based approaches: we do not modify model outputs directly and require only black-box access, while naturally supporting multi-bit message embedding and generalizing across diverse languages and domains. We instantiate the framework using Sparse Autoencoders as deterministic feature extractors and provide theoretical worst-case analysis relating watermark accuracy to computational budget. Experiments across 4 datasets demonstrate strong watermarking performance on English, Chinese, and code while preserving text quality. SAEMark establishes a new paradigm for scalable, quality-preserving watermarks that work seamlessly with closed-source LLMs across languages and domains.
@@ 1 Introduction
Figure 1: An overview of SAEMark.
Large language models (LLMs) have revolutionized text generation across domains, from creative writing to code synthesis [6, 10]. However, their ability to produce human-quality text at scale raises serious concerns about misinformation, copyright infringement, and content laundering. As these models become ubiquitous, reliably attributing AI-generated content becomes critical for accountability and trust.
Watermarking—embedding detectable signatures into generated text—offers a promising solution. They must preserve text quality while enabling reliable detection, operate across languages and domains, and scale to distinguish between many users or sources. Most critically, they must work with real-world deployment constraints where model providers offer only API access without exposing internal parameters. The challenge becomes even more complex for multi-bit watermarking. Beyond simply detecting AI-generated text, the goal is to encode and recover a specific message $m\in\{0,1\}^{b}$—such as a user identifier for personalized attribution. This enables answering not just “is this AI-generated?” but “which specific user or system generated this text?” Such fine-grained attribution is essential for large-scale deployment where accountability matters.
Existing watermarking methods struggle with these requirements. Token-level approaches like KGW [16] and EXP [1] require direct access to model logits, excluding API-based deployment, and can degrade text quality through probability manipulation. Syntactic methods [11] fail to generalize across languages, while specialized approaches [18] work well in narrow domains but break down when applied more broadly. Even recent black-box methods [4, 7] rely on surface-level statistics or require auxiliary models, limiting their robustness and scalability.
@@
This approach works by operating on meaningful units of text—sentences for natural language, functions for code. For each unit, we extract deterministic features that capture semantic properties, compute a scalar statistic, and normalize it to behave predictably across different texts. Using the watermark key, we derive target values for each position. During generation, we sample multiple candidates from the LLM and select the one whose feature statistic is closest to the target, ensuring the final sequence encodes the desired message. The elegance lies in what we don’t change: no model weights, no logit manipulation, no token modifications. Every selected text segment is a natural LLM output, preserving quality while enabling attribution. The approach works with any LLM through API calls, generalizes across languages and domains, and provides theoretical guarantees on watermark success that scale predictably with computational budget.
@@
Our contributions span theory and practice. We develop a general framework for watermarking through feature-guided selection that works with any feature extractor and LLM. We provide theoretical guarantees that explain how SAEMark accuracy scales with compute budget, independent of feature extractors. Finally, we demonstrate a practical instantiation using SAE as feature extractor that achieves superior accuracy and text quality across languages and domains, encoding more information than existing approaches.
2 Related Work
@@
LLM watermarking is a technique to embed special patterns into the output of LLMs, and has traditionally been used to identify LLM generated text from human-written text [14]. Different from post-hoc detection methods [31] that analyze statistical patterns in existing text, language model watermarking aims to embed detectable signatures during generation [16]. These methods compromise generation quality through direct manipulation of token probabilities [16] or syntactic modifications [2]. The challenge of language and domain generalization remains largely unaddressed, with current techniques primarily optimized for English and struggling with multilingual content or specialized domains like code [18]. Notably, PersonaMark [33] represents early attempts at personalized watermarking, but its reliance on English-specific syntactic patterns and closed-source implementation makes scalability and cross-lingual capability difficult to verify. Recently, more multi-bit watermarking methods have been proposed to embed multiple bits of information into generated text [17, 32, 29, 9, 30, 22], primarily by extending single-bit watermarking that manipulates logits during generation; these methods inherit the limitations of single-bit designs.
@@
Complementary black-box watermarks avoid white-box logit access by using post-hoc selection or rewriting [7]. However, they typically operate on statistics or introduce auxiliary model dependencies and do not directly address multi-bit message embedding at scale. Our framework differs by performing inference-time selection among naturally generated candidates using deterministic feature statistics. This enables extractor-agnostic analysis and multilingual, domain-agnostic multi-bit watermarking without modifying model logits.
@@
Sparse Autoencoders (SAEs) are pre-trained interpretability tools that decompose LLM activations into human-understandable features [5]. For a given base model $\mathcal{M}$ and layer $l$, an SAE processes hidden states $\mathbf{h}_{t}$ at position $t$ as:
@@
This training produces features that correspond to interpretable concepts [5, 24] like “Function definitions” or “Concept related to color blue” [19]. Our watermarking leverages key properties of pre-trained SAEs: multilingual activation allows the same features to fire for equivalent concepts across languages deterministically. sparsity enables efficient analysis through few active features per token. These properties support language-agnostic statistics via masked feature aggregations:
@@
where $\mathbf{m}$ filters background features that fire ubiquitously regardless of content (e.g., punctuation). The summary $\phi(\mathbf{y})$ provides a deterministic statistic used by our watermarking procedures (Sec 3).
3 Methodology
@@ 3.1 Task Definition
Given a base LLM $\mathcal{M}$, watermark key $k\in\mathcal{K}$, input $\mathbf{x}$, and message $m\in\{0,1\}^{b}$, the algorithm produces $\mathbf{y}$ by post-hoc selection over $\mathcal{M}$’s outputs (black-box/API-compatible; no access to logits or parameters; cf. black-box watermarking [4, 7]):
@@
Threat model and scope: The scheme targets three properties: key privacy (deriving $k$ from watermarked outputs is hard), verifier-held detectability (any party holding $k$ can verify), and collusion resistance (multiple keys should not facilitate forgery). Our focus is attribution without storing LLM generated text. This work does not claim cryptographic unforgeability when keys are known; preventing adversarial forgeries is an important direction for security-focused follow-ups.
@@ 3.2 General Framework for Feature-based Watermarking
Input: Prompt $c$, key $k$, LLM $G$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, units $M$, attempts $K$, candidates NN
Output: Watermarked text $x^{*}$
for $attempt\leftarrow 1$ to $K$ do
  for $i\leftarrow 1$ to $M$ do
     $\mathcal{X}\leftarrow\text{GenerateCandidates}(G,x^{*},N)$
  end for
 if CheckAlignment($\{\tau_{i}\},\{z_{i}\}$) then
    return $x^{*}$
  end if
end for
return $x^{*}$
Algorithm 1 Watermark generation
Input: Text $x$, candidate keys $\mathcal{K}$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, alignment thresholds, significance $\alpha$
Output: Detection result $d\in\mathcal{K}\cup\{\emptyset\}$
$\mathcal{D}\leftarrow\emptyset$
foreach $k_{i}\in\mathcal{K}$ do
  $\{\tau_{j}\}\leftarrow\text{TargetsFromKey}(k_{i},|\{z_{j}\}|)$
  if CheckAlignment($\{\tau_{j}\},\{z_{j}\}$) then
     $t,p\leftarrow\text{StudentTTest}(\{z_{j}\},\{\tau_{j}\})$
     if $t>t_{\alpha/2}\land p<\alpha$ then
        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(k_{i},t)\}$
     end if
  end if
end foreach
Algorithm 2 Watermark detection
Figure 2: Pseudocode for SAEMark: generation and detection.
@@
where $M$ denotes the number of textual units in the sequence and $O_{\min}=0.95$ ensures that at least 95% of target values fall within the observed range. These two filters aim to eliminate spurious matches: the range similarity filter prevents matching sequences with fundamentally different statistical properties, while the overlap rate filter ensures meaningful correspondence between target and observed values. Only after passing both alignment checks do we apply Student’s t-test for statistical significance. The key with the highest significance score is returned if it passes the threshold; otherwise, we classify the text as unwatermarked.
@@ 3.3 Theoretical Analysis and Guarantees
Worst-case analysis and bounds.
@@
where $\Phi$ denotes the standard normal CDF and $p_{\min}=\Phi(2(1+k)+k\mu/\sigma)-\Phi(2(1-k)-k\mu/\sigma)$.
@@ 3.4 Sparse Autoencoder Instantiation
What concrete feature extractor should we use? We need statistics that are deterministic, semantically meaningful, and statistically regular. Sparse autoencoders—interpretability tools designed to understand language model internals—provide an ideal solution. They decompose language representations into interpretable semantic components (e.g. “technical writing,” “math symbols”) that exhibit distinctly different activation patterns across generations. By applying the SAE to a separate “anchor” model, our approach remains compatible with any target language model, including API services, while extracting the discriminative yet predictable statistics our framework requires.
@@
Two practical optimizations bridge the gap between our theoretical framework and robust empirical performance. First, the CheckAlignment algorithm’s eliminate spurious statistical matches that would otherwise compromise detection accuracy. Second, background feature masking ensures FCS calculations focus on discriminative semantic patterns rather than ubiquitous surface features. We precompute a mask excluding "background" SAE features, like those related to punctuation or basic grammar, to focus on discriminative semantic patterns. With empirically observed parameters $\mu=0.142$, $\sigma=0.029$, and tolerance $\epsilon=0.1$, our bounds yield concrete success probabilities: $N=50$ achieves $99\%$ per-unit success, $N=20$ maintains $85.32\%$, and $N=10$ achieves $61\%$. Since each generation involves multiple units, overall success rates significantly exceed these per-unit bounds. Modern inference engines support parallel generation of $N$ candidates simultaneously, making the approach practically efficient despite these extra compute overhead. We have extensive ablations and empirical results in the following experiments.
@@ 4.1 Experimental Setup
C4 [23] LCSTS [12] MBPP [3] PandaLM [26]
@@
Task Type Completion Summarization Code Generation Instruction Following
@@
We evaluate on 4 diverse datasets as shown in Table 1. Following common practice in prior work, we report Accuracy, Recall, F1 at 1% FPR. For text quality, we report win-rates of pairwise comparison on PandaLM judged by GPT-4o in our main results, and average pointwise scores on BIGGen-Bench judged by their officially released judge model as an alternative text-quality experiment. We use implementation for single-bit watermarks from MarkLLM [21] toolkit and Waterfall [17] as it’s the current best open-source training-free multi-bit watermark similar to our setting. Full details of baselines in Appendix C.
@@ 4.2 Watermarking Accuracy and Text Quality
Method C4 (English, [23])     LCSTS (Chinese, [12]) MBPP (Code, [3])     PandaLM (Instruction,  [26])
@@
KGW [16] 99.2 99.6 99.2 99.1 98.8 99.1 65.4 31.9 48.0 41.5 89.9 80.4 88.8
EXP [1] 99.5 99.6 99.5 99.3 99.4 99.3 57.8 16.7 28.4 23.2 79.3 59.4 74.2
UPV [20] 86.0 72.0 83.7 90.5 91.0 90.5 51.6 3.1 6.0 36.0 54.0 8.0 14.8
Unigram [34] 98.8 98.6 98.8 98.2 97.0 98.2 65.4 31.9 48.0 35.3 53.3 7.2 13.4
DIP [27] 96.0 92.6 95.9 97.7 96.2 97.7 60.7 22.6 36.5 36.5 81.5 63.8 77.5
Unbiased [13] 96.7 94.4 96.6 97.8 96.4 97.8 64.0 29.2 44.8 40.2 74.3 49.3 65.7
SynthID [8] 98.2 97.2 98.2 97.6 96.2 97.6 62.5 26.1 41.0 36.0 81.2 63.0 77.0
SWEET [18] 99.6 99.6 99.6 50.0 0.0 0.0 72.4 45.9 62.4 47.2 87.7 76.8 86.2
@@
Beyond accuracy, Table 2 shows SAEMark achieves the highest quality score (67.6%) on PandaLM as judged by GPT-4o pairwise comparisons. To study how this generalizes across different backbone LLMs, we conduct additional evaluation on BIGGen-Bench comparing against watermarked baselines and unwatermarked text.
Table 3: Text quality evaluation on BIGGen-Bench. Scores are on a 5-point Likert scale (higher is better) [15].
@@
Table 3 confirms SAEMark achieves the highest quality among watermarks across three backbone LLMs. This quality advantage stems from our design: rather than manipulating logits or applying external rewriting to obtain watermarked text, we simply run post-hoc selection among naturally generated candidates, ensuring text quality stays bounded by the its own capabilities.